Suppose the distribution of distribution of the number of days the typical college student consumes alcohol per year has a mean of μ = 90 and σ = 7.2. If a student consumes alcohol less often than 90% of his or her peers, how often does he or she drink?
Answer by Boreal(15235)

the 10th percentile has z= -1.282
z=(x-mean)/sd
so-1.282=(x-90)/7.2
-9.23=x-90
x=80.77 days or 81 days a year.